OpenStudy (anonymous):

5/6b -1/3 =3/2b what does b = ?

OpenStudy (anonymous):

b = -4

OpenStudy (anonymous):

(5/6b)-(1/3)=(3/2b) (5/6b)-(3/2b)=1/3 (1/b)*((5/6)-(3/2))=1/3 on further simplification 3*(-4)/6=b b=-2